40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 27 15:03:02 2016, sisyphus, Metaphysics, SUS, Broken wire Unknown.jpeg
    Reply  Wed Jul 27 16:56:47 2016, rana, Metaphysics, SUS, Broken wire 
    Reply  Thu Jul 28 08:42:36 2016, Steve, Update, SUS, wire clamp e12342.jpg
    Reply  Fri Jul 29 15:17:23 2016, Steve, Update, SUS, wire and clamping test 43micr242grs.jpg
       Reply  Fri Jul 29 20:12:54 2016, gautam, Update, SUS, third time - no luck 
          Reply  Mon Aug 1 14:09:20 2016, Steve, Update, SUS, third time - no luck 
             Reply  Tue Aug 9 16:19:27 2016, Steve, Update, SUS, torque driver for wire clamp 
                Reply  Wed Aug 10 15:34:24 2016, Steve, Update, SUS, 6 in-lbs torque driver for wire clamp screw 
          Reply  Wed Aug 3 00:15:39 2016, gautam, Update, SUS, ETMX suspended 
             Reply  Wed Aug 3 09:26:54 2016, Lydia, Update, SUS, ETMX suspended: photos 8x
             Reply  Thu Aug 4 03:14:39 2016, gautam, Update, SUS, ETMX suspended 
                Reply  Fri Aug 5 02:17:10 2016, gautam, Update, SUS, ETMX Ruby Wire Standoff Ready for Gluing 8x
                   Reply  Fri Aug 5 04:43:09 2016, Koji, Update, SUS, ETMX Ruby Wire Standoff Ready for Gluing 
                   Reply  Fri Aug 5 18:21:28 2016, gautam, Update, SUS, ETMX Ruby Wire Standoff GLUED 11x
                      Reply  Sun Aug 7 14:53:39 2016, ericq , Update, SUS, ETMX Standoff gluing was successful ETMX_resonances.pdfETMX_SUSspectra.zip
                         Reply  Tue Aug 9 00:44:43 2016, gautam , Update, SUS, ETMY patch-up 
                            Reply  Tue Aug 9 15:27:57 2016, gautam , Update, SUS, ETMY patch-up IMG_2996.JPGIMG_2997.JPG
                               Reply  Wed Aug 10 03:08:03 2016, gautam , Update, SUS, ETMY patch-up 
                                  Reply  Wed Aug 10 23:45:03 2016, gautam , Update, SUS, ETMY re-suspended 7x
                                     Reply  Thu Aug 11 00:20:41 2016, Koji, Update, SUS, ETMY re-suspended 
                                        Reply  Thu Aug 11 11:56:40 2016, gautam, Update, SUS, ETMY re-suspended pitchBalancingWithBias.PDF
                                           Reply  Mon Aug 15 18:28:15 2016, gautam, Update, SUS, Air-bake preparation IMG_6816.JPG
                                              Reply  Tue Aug 16 08:04:49 2016, Steve, Update, SUS, ETMY had Aluminum bridge AlBridge.jpg
                                              Reply  Tue Aug 16 21:54:27 2016, gautam, Update, SUS, Air-bake - IN PROGRESS IMG_3005.JPG
                                                 Reply  Wed Aug 17 08:47:43 2016, ericq, Update, SUS, Air-bake finished 
                                                    Reply  Wed Aug 17 23:00:57 2016, gautam, Update, SUS, Air-bake of towers IMG_3006.JPG
                                                       Reply  Thu Aug 18 08:17:16 2016, Steve, Update, SUS, wire clamping preparation 800P.jpg
                                                       Reply  Thu Aug 18 14:14:20 2016, gautam, Update, SUS, Air-bake of towers - finished 
                                                          Reply  Fri Aug 19 22:51:12 2016, gautam, Update, SUS, ETMs first-contacted 
                                                             Reply  Sun Aug 21 16:23:05 2016, gautam, Update, SUS, ETMX suspended IMG_3018.JPG
                                                                Reply  Mon Aug 22 13:06:11 2016, gautam, Update, SUS, ETMY suspended IMG_3019.JPG
                                                                   Reply  Mon Aug 22 18:04:24 2016, gautam, Update, SUS, ETMY OSEMs inserted 
                                                                      Reply  Tue Aug 23 19:35:38 2016, gautam, Update, SUS, ETMY back in IFO 10x
                                                                         Reply  Thu Aug 25 21:11:43 2016, gautam, Update, SUS, ETMY back in IFO ETMY_BounceSpectra_25Aug2016.pdf
                                                                            Reply  Fri Aug 26 11:35:44 2016, gautam, Update, SUS, ETMY UL sensor problematic  ETMY_BounceSpectra_26Aug2016.pdf41.png19.png
                                                                               Reply  Fri Aug 26 14:05:38 2016, Steve, Update, SUS, ETMY UL sensor problematic  ETMY-ULhistory.png
                                                                               Reply  Fri Aug 26 14:09:46 2016, ericq, Update, SUS, ETMY UL sensor problematic  
                                                                                  Reply  Sun Aug 28 21:40:11 2016, gautam, Update, SUS, ETMY UL sensor problematic  ETMY_UL_problematic.png
                                                                                     Reply  Mon Aug 29 09:37:05 2016, gautam, Update, SUS, ETMY UL sensor problematic  satelliteBox.png
                                                                                        Reply  Mon Aug 29 14:42:06 2016, gautam, Update, SUS, ETMY Satellite box diagnostics IMG_6840.JPGIMG_6841.JPG
                                                                            Reply  Fri Aug 26 17:48:14 2016, gautam, Update, SUS, bounce mode coupling reduction ETMY_BounceSpectra_26Aug2016_1.pdf
                                                                               Reply  Mon Aug 29 21:53:14 2016, Lydia, Update, SUS, ETMY bounce mode coupling IMG_3051.JPG42.png
                                                                                  Reply  Tue Aug 30 20:26:36 2016, gautam, Update, SUS, ITMY installed IMG_3052.JPGSRM_sensor_level_comp.pngETMY_BounceSpectra_30Aug2016.pdf
                                                                                     Reply  Wed Aug 31 18:07:43 2016, Johannes, Update, SUS, ITMY free swinging 
                                                                                        Reply  Wed Aug 31 22:09:18 2016, gautam, Update, SUS, Y arm locked to Green  
                                                                                           Reply  Thu Sep 1 08:30:24 2016, Steve, Update, SUS, ITMY_UL is sick ITMY_SENSOR_UL.pngspareSatAmp.jpg
                                                                                              Reply  Tue Sep 6 20:30:56 2016, rana, Update, SUS, ITMY_UL is sick 
                                                                                                 Reply  Tue Sep 13 11:00:59 2016, Lydia, Update, SUS, ETMY UL glitch returned 
                                                                                           Reply  Thu Sep 1 19:59:22 2016, Johannes, Update, SUS, IR mode flashes in Y arm 
                                                                                              Reply  Fri Sep 2 21:09:08 2016, gautam, Update, SUS, Y arm locked in air 03.png56.png
                                                                                                 Reply  Tue Sep 6 20:45:14 2016, gautam, Update, SUS, X arm test masses back in chamber IMG_3137.JPGIMG_3143.JPGIMG_3142.JPGIMG_3148.JPG
                                                                                                    Reply  Thu Sep 8 22:12:36 2016, gautam, Update, SUS, X arm in place, locked to green, IR flashes visible 
                                                                                                       Reply  Fri Sep 9 17:50:02 2016, gautam, Update, SUS, Heavy doors on BS-PRM, ETMY chambers 
                                                                                                          Reply  Sat Sep 10 14:26:01 2016, rana, Update, General, PRM SRM alignment Osamu.JPG
                                                                                                             Reply  Mon Sep 12 17:15:22 2016, Johannes, Update, General, PRM SRM alignment IMG_3273.JPGIMG_3275.JPG
                                                                                        Reply  Thu Sep 1 19:18:14 2016, Johannes, Update, SUS, ITMX and ETMX preemptive table leveling 8x
                                                                                           Reply  Fri Sep 2 21:15:53 2016, gautam, Update, SUS, ITMX and ETMX preemptive table leveling 
                                                                                     Reply  Thu Sep 1 15:43:16 2016, rana, Update, SUS, OSEM adjustments 
                                                                                        Reply  Mon Sep 5 19:57:24 2016, Osamu, Update, SUS, OSEM adjustments 
Message ID: 12434     Entry time: Tue Aug 23 19:35:38 2016     In reply to: 12430     Reply to this: 12444
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY back in IFO 

[johannes, gautam]

Summary: Today we moved the suspended ETMY optic back into the chamber from the cleanroom. Once in the chamber, we positioned the optic using the stops that marked the previous position of the optic. We then shortened the arm length by 19mm (in order to match the X and Y arm lengths. The F.C. coat on the HR face was removed prior to the final placement of the optic. We then adjusted the OSEM positions in their holders to get the sensor outputs to half their maximum value.

We did not get to check where the input beam hits the optic or see if the pitch balance of the optic is such that the reflected beam makes it back to the ITM. The plan for tomorrow is to do this. 


Part 1: Cleanroom work

  • We worked a little more on trying to adjust the rotational position of the OSEM coils in order to minimize the coupling of the bounce mode into the sensor signals. 
  • We had limited success in this regard. After about an hour, we concluded that it made more sense to do this in the chamber itself. For one thing, the drive electronics for the Y end are different (in the cleanroom, we are using the X end electronics, satellite box etc.).
  • We adjusted the position of the OSEMs till the sensor output readout was half the open value as best as we could. We also made sure that the wire was in the groove on both sides and that the magnets were well centered in the vertical direction relative to the OSEM coils and that there was no danger of knocking any magnets off (see attached pictures).
  • We then engaged all the EQ stops, and transferred the suspension cage to a cart (topped with Al foil, wiped clean) for transportation to the Y-end (with OSEMs left in).

Part 2: Transportation of optic

  • Nothing special here, just took great care while going over bumps near doors between the cleanroom and the IFO, and along the Y-arm itself.
  • Definitely a 2 man job - one person can lift a pair of wheels over any bumps while the other can make sure there is no danger of the cage toppling over. 

Part 3: Chamber work

  • PSL shutter was closed for this part of work. Earlier today, I found that C1SUSAUX had failed yet again (why are all the slow computers dying more often nowadays?!). I restarted the slow machine, and locked the mode cleaner. The alignment hadn't drifted so much from when EricQ had last aligned the IMC, and with only minimal tweaking, I was able to lock the IMC and see a beam on the REFL camera.
  • First, I transferred the suspension cage onto the edge of the table inside the chamber. Care was taken not to accidentally place the cage onto the trailing OSEM wires.
  • There were some specks of dust on the barrel of the optic, and also the cage. These were removed with clean wipes and isopropanol.
  • I judged that it would be too precarious to remove the F.C. with the optic in its final desired position. So we decided to take the coat off with the optic at the edge of the table. The central part of the HR face looks pretty clean. Even though the whole HR face was cleaned with F.C., the part that was left uncovered prior to putting the optic back into the chamber has a few specks of dust on it (see attachments). These could not be removed just by blowing ionized air. I was hesitant to drag wipe the optic, so I left things as is. In any case, the optic as a whole is MUCH cleaner (to my eye at least) than prior to the cleaning. 
  • Conveniently, the stops marking the previous position of the optic were on the far side and back.
  • Since we wanted to shorten the Y arm length by 2 cm, we placed a clean steel ruler of width 19mm in front of the rear stop (see attached pictures). I then moved the cage back along the side stop till I hit the ruler.
  • I then clamped the optic down, removed the spacing ruler, and re-adjusted the position of the rear stop to mark the new position of ETMY.
  • We were concerned that the change of position of the cage on the table affected the leveling. Checking with a clean spirit level, we found evidence of a slight tilt in the direction towards the vertex of the IFO, as expected from the way the ETMY cage was moved. To compensate for this, I moved one of the counterweight masses (see attachments) till the spirit level showed the table to be level (to its resolution) in two perpendicular directions
  • We then plugged in the OSEMs into the DB25 connectors on the table. We found that the Y-end electronics were giving different readouts from what we had been seeing in the cleanroom with the X end electronics (not surprising I guess). We resolved to pull out all the OSEMs, check their maximum sensor output values, and re-insert them till the sensor output was half this maximum as best as we could. NOTE TO SELF: UPDATE THE WIKI PAGE!
  • We turned on the damping, and found that the exisiting input matrix performs fairly well.
  • We took a quick look at the spectra of the sensor outputs - interestingly, with the suspension on the seismic stacks inside the chamber, the 16.4 Hz bounce mode peak showed up clearly (these were totally absent in the cleanroom). I did not attempt any fine rotation of the OSEMs in the holders (it is not even clear to me how good/bad the present configuration is) because I reasoned we first need to apply a pitch bias to get the beam back to the ITMY chamber and then re-adjust the OSEM coils. The bounce mode decoupling will be the last step. 
  • For tonight, we decided to leave the optic freely swinging (with EQ stops close by) so that tomorrow, we can look at the offline spectra of sensor outputs and if necessary, re-diagonalize the suspension. 
  • After checking nothing unwanted was left behind in the chamber, we closed it up for tonight.

Plan for tomorrow:

  • Pitch balancing check (by looking at reflected beam at ITMY)
  • Re-adjust OSEMs on ETMY, minimize bounce mode coupling into sensor outputs
  • Make Y arm cavity by re-positioning ITMY

Attachments:

Attachment #1: Wire is in groove in side without OSEM

Attachment #2: Wire is in groove in side with OSEM (picture taken with OSEM coil removed)

Attachment #3: UL magent relative to OSEM coil

Attachment #4: LL magent relative to OSEM coil

Attachment #5: LR magnet relative to OSEM coil

Attachment #6: UR magnet relative to OSEM coil

Attachment #7: Side magnet relative to OSEM coil

Attachment #8: ETMY HR face with F.C. film removed. Non-covered part isn't super clean, but the covered part itself does not have any large specks of dust visible.

Attachment #9: Scheme adopted to shorten Y arm length by 19mm.

Attachment #10: Current situation inside EY chamber. Counterweight that was moved to balance the table is indicated.

 

Attachment 1: IMG_3025.JPG  6.427 MB  Uploaded Tue Aug 23 22:39:09 2016  | Hide | Hide all
IMG_3025.JPG
Attachment 2: IMG_3035.JPG  6.360 MB  Uploaded Tue Aug 23 22:40:14 2016  | Hide | Hide all
IMG_3035.JPG
Attachment 3: IMG_3030.JPG  6.014 MB  Uploaded Tue Aug 23 22:40:59 2016  | Hide | Hide all
IMG_3030.JPG
Attachment 4: IMG_3029.JPG  6.014 MB  Uploaded Tue Aug 23 22:41:30 2016  | Hide | Hide all
IMG_3029.JPG
Attachment 5: IMG_3028.JPG  6.230 MB  Uploaded Tue Aug 23 22:42:31 2016  | Hide | Hide all
IMG_3028.JPG
Attachment 6: IMG_3027.JPG  5.700 MB  Uploaded Tue Aug 23 22:43:25 2016  | Hide | Hide all
IMG_3027.JPG
Attachment 7: IMG_3026.JPG  5.828 MB  Uploaded Tue Aug 23 22:43:40 2016  | Hide | Hide all
IMG_3026.JPG
Attachment 8: IMG_3036.JPG  5.212 MB  Uploaded Tue Aug 23 22:45:25 2016  | Hide | Hide all
IMG_3036.JPG
Attachment 9: IMG_3038.JPG  5.397 MB  Uploaded Tue Aug 23 22:45:46 2016  | Hide | Hide all
IMG_3038.JPG
Attachment 10: IMG_3045.JPG  5.203 MB  Uploaded Tue Aug 23 22:49:45 2016  | Hide | Hide all
IMG_3045.JPG
ELOG V3.1.3-